OpenStudy (anonymous):

Identify the solutions of 3x^2 − 13x + 4 = 0. x = −1 over 3, x = −4 x = 1 over 3, x = 4 x = −4 over 3, x = −1 x = 4 over 3, x = 1

OpenStudy (alexandervonhumboldt2):

do ou know quadratic formual

OpenStudy (anonymous):

No what is that?

OpenStudy (anonymous):

factor

OpenStudy (wolf1728):

x = [-b +-square root (b^2 -4ac) ]/ 2a

OpenStudy (anonymous):

what does that mean??

OpenStudy (wolf1728):

In your equation a=3 b= -13 and c =4

OpenStudy (wolf1728):

So x = [13 +-sq root(169 -4*3*4)]/2*3 Can you calculate it from there?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (wolf1728):

Okay and just to check your answer you can go here: http://www.1728.org/quadratc.htm

OpenStudy (anonymous):

C right?

OpenStudy (misty1212):

factor

OpenStudy (anonymous):

c?

OpenStudy (misty1212):

\[(3x-1)(x-4)=0\] and go from there

OpenStudy (misty1212):

no

OpenStudy (misty1212):

\[x-4=0\\ x=4\]

OpenStudy (misty1212):

\[3x-1=0\\ 3x=1\\ x=\frac{1}{3}\]

OpenStudy (anonymous):

Its B ya?